OpenStudy (anonymous):

Please Help Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (unklerhaukus):

what is the the difference between the equations ?

OpenStudy (anonymous):

-2x and =4x 1 and -1

OpenStudy (unklerhaukus):

what is -2x - 4x = ? what is 1 - (-1) = ??

OpenStudy (anonymous):

-6x and 2

OpenStudy (unklerhaukus):

whops i meant -2x - (-4x) - ?

OpenStudy (unklerhaukus):

-2x + y = 1 -4x + y = -1 the difference between the left hand sides of the equations are (-2x + y) - (-4x + y) = -2x + y + 4x - y = -2x + 4x = ?

OpenStudy (anonymous):

okay 6x

OpenStudy (unklerhaukus):

= (-2 + 4)x =?

OpenStudy (unklerhaukus):

you already got the difference between the right hand sides 1 - (-1) = 1+1 = 2

OpenStudy (anonymous):

Oh, didn't see the negative lol = (-2+4)x =2x

OpenStudy (unklerhaukus):

so the difference between the two equation -2x + y = 1 -4x + y = -1 becomes -2x + y - (-4x+y) = 1- (-1) 2x = 2

OpenStudy (unklerhaukus):

solve this for x

OpenStudy (unklerhaukus):

then sub this value into either -2x + y = 1 OR -4x + y = -1 and solve for y the solution is (x, y)

OpenStudy (anonymous):

Okay

OpenStudy (unklerhaukus):

what'd you get for x?

OpenStudy (anonymous):

I'm confused sorry

OpenStudy (unklerhaukus):

can you see how we got from -2x + y - (-4x+y) = 1- (-1) to 2x = 2 ?

OpenStudy (anonymous):

Nope

OpenStudy (unklerhaukus):

do you know where -2x + y - (-4x+y) = 1- (-1) came from?

OpenStudy (anonymous):

Yeah I know where that was from

OpenStudy (unklerhaukus):

well the right hand side -2x + y - (-4x+y) becomes -2x + y + 4x - y -2x + 4x +y - y -2x + 4x 2x

OpenStudy (unklerhaukus):

and the right hand side = 1- (-1) = 1 + 1 = 2

OpenStudy (unklerhaukus):

so we have found the difference between the equations to be 2x = x

OpenStudy (anonymous):

okay I still don't get how I find x from -2x + y - (-4x+y) = 1- (-1) 2x = 2

OpenStudy (unklerhaukus):

2x = 2 divide both sides of the equation by 2 2x /2 = 2/2 x =

OpenStudy (anonymous):

Oh I was confused because I was looking at the whole thing x= 1

OpenStudy (unklerhaukus):

good! you are almost there now,

OpenStudy (unklerhaukus):

now sub x= 1 into either one of the original equations -2x + y = 1 OR -4x + y = -1 and solve for y

OpenStudy (anonymous):

(-2x1) + y = 1 -2+ y= 1

OpenStudy (unklerhaukus):

that's right ....

OpenStudy (unklerhaukus):

y = ???

OpenStudy (anonymous):

i think y=3

OpenStudy (unklerhaukus):

yes!

OpenStudy (unklerhaukus):

so you have found x = 1, y= 3 ie (x, y) = (1, 3)

OpenStudy (anonymous):

Cool, I missed out on class today but I still had homework. So I needed a little help

OpenStudy (unklerhaukus):

if you plug in these values into the original equations -2x + y = 1 OR -4x + y = -1 you'll get -2(1) + 3 = 1 OR -4(1) +3 = -1 -2 + 3 = 1 OR -4 +3 = -1 1 = 1 OR -1 = -1 which are both always true so you know your solution is correct

OpenStudy (anonymous):

Okay

OpenStudy (unklerhaukus):

do you have another one like this?

OpenStudy (anonymous):

I have a couple

OpenStudy (anonymous):

But now I don't have time to solve more.

OpenStudy (unklerhaukus):

we found the difference of the equation for this one because it made the y's cancel out , sometime you have to add the equations, or something more complex

OpenStudy (anonymous):

Okay. Thank you sooo much. I have to leave. I'll post another question if I have problems

OpenStudy (unklerhaukus):

ok \[\ddot\smile\]